OpenStudy (anonymous):

how to factor 5x^2-32x-21=0

hero (hero):

It's factorable, but not in the way that you're used to doing it

OpenStudy (anonymous):

Solve for the roots of the quadratic.

OpenStudy (anonymous):

ok and how exactly you do that

hero (hero):

I'll post the steps

OpenStudy (anonymous):

Quadratic equations take the form: \[\Large ax{^2} + bx + c = 0\] Use the quadratic formula to find its roots: \[\Large \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\]

OpenStudy (anonymous):

well I need to use a factoring way not a quadratic way, I already know how to do it by the qudratic form

hero (hero):

5x^2-32x-21=0 5x^2+3x - 35x - 21 = 0 x(5x+3)-7(5x+3) = 0 (5x+3)(x-7) = 0 x = 7, -3/5

OpenStudy (anonymous):

I don't get the second line

OpenStudy (anonymous):

-32 = -35+3

hero (hero):

Split the middle term into 3 and -35 adds to get -32 multiplies to get -105

OpenStudy (anonymous):

ok

hero (hero):

-105 is what you get when you multiply a*c or 5*-21

OpenStudy (anonymous):

I don't get how you did the 3rd line